OpenStudy (anonymous):

Find y' by implicit differentiation: 1/x + 1/y = 1

sam (.sam.):

\[\frac{1}{x}+\frac{1}{y}=1\] \[x^{-1}+y^{-1}=1\] -------------------------------------- Differentiate, \[-x^{-2}-y^{-2}(y')=0\] \[-y^{-2}(y')=x^{-2}\] \[(y')=\frac{x^{-2}}{-y^{-2}}\] \[y'=-\frac{y^2}{x^2}\]

OpenStudy (anonymous):

Okay so now: Solve the equation explicitly for y and differentiate to get y' in terms of x.